OpenStudy (anonymous):

Can someone help me with question 4 and 5 please. Will be much appreaciated😄

Directrix (directrix):

Please post questions 4 and 5. Thanks.

OpenStudy (anonymous):

Didn't u get the attachment

OpenStudy (anonymous):

Okay I will type it

OpenStudy (anonymous):

Question 4: To the nearest centimetre find how far a tent peg is away from a tent pole if the 3m guy rope is pegged so that it makes an angle of 50 degrees with the ground.

OpenStudy (anonymous):

Question 5: The railway to Angleton rises 35cm for every 50m of track. Find the angle at which the track is inclined to the horizontal, to the nearest minute.

Directrix (directrix):

>>>Didn't u get the attachment No

OpenStudy (anonymous):

Did you get a picture?

Directrix (directrix):

I did not get the picture of the tent until now. Use right triangle trig. The cosine of 50 degrees = x/ 3 where x is the distance of the peg from the pole. cos(50) = x/3 x = 3* cos(50) x = ? @132231

Directrix (directrix):

Review the right triangle trig ratios. See attachment.

Directrix (directrix):

Which trig ratio should be used here?